Your answer is incorrect. Try again.
An analysis of the transactions made by Skysong, Inc., a certified public accounting firm, for the month of August is shown below.
Assets = Liabilities + Stockholders’ Equity


Cash


+

Accounts
Receivable


+

Supplies


+

Equipment

Accounts
Payable

Common
Stock

Retained Earnings
= + +

Revenues

Expenses

Dividends

1. $16,000 $16,000
2. –2,500 $5,500 $3,000
3. –700 $700
4. 5,700 $3,800 $9,500 Service Revenue
5. –1,300 –1,300
6. –2,000 -$2,000
7. –600 –$600 Rent Expense
8. 350 –350
9. –4,100 –4,100 Salaries/Wages Expense
10. 300 –300 Utilities Expense

(b) Determine how much stockholders’ equity increased for the month.
Increase in stockholders' equity $Enter the increase in stockholders' equity in dollars

Entry field with incorrect answer


(c) Compute the amount of net income for the month.
Net income $Enter the net income in dollars

Entry field with incorrect answer

Click if you would like to Show Work for this question:

Open Show Work

LINK TO TEXT

LINK TO TEXT

Question Attempts: 3 of 15 used

SAVE FOR LATER

SUBMIT ANSWER
Answer #1

a) Increase in Stockholder's equity = Issue common Stock+Net income-Dividend = 16000+4500-2000 = 18500

b) Net income = 9500-600-4100-300 = 4500